OpenStudy (anonymous):

Please help me precalc? A wheel with radius 5 cm is being pushed up a ramp at a rate of 7 cm per second. The ramp is 760 cm long, and 260 cm tall at the end 1)Traveling at 7 cm/sec, in one second, how much does the height of the center of wheel increase? 2)Write an equation for the height of the center of the wheel as a function of t, time in seconds. center height= 3)Write an equation for the height of the point P above the ground as a function of t, time in seconds. h(t) =

OpenStudy (anonymous):

I calculated this so far: the wheel travels 2*pi*5=31.41592654cm in one rotation. If the wheel travels 2*pi*5=31.41592654cm/sec it will take 31.41592654/7=4.488seconds to complete one rotation. The initial height is 5cm??...

OpenStudy (anonymous):